# NS-FPN: Improving Infrared Small Target Detection and Segmentation from Noise Suppression Perspective

### v1

@@ 4.1 Datasets and Evaluation Metrics
Table 3: LFP at different scale layers are performed on the IRSTD-1k dataset.
@@
4.2 Implementation Details
All experiments are conducted on NVIDIA GeForce RTX 4090 GPUs. We integrate our NS-FPN into MSHNet (Liu et al. 2024) for segmentation and into YOLOv8n-p2 for detection. The Adagrad optimizer (Duchi, Hazan, and Singer 2011) is used with an initial learning rate of 0.05. The models are trained for 500 epochs with a batch size of 16. For segmentation tasks, input images are resized to 256 and then randomly cropped to 224 during training. For detection tasks, input images are uniformly resized to 640.
@@ 4.2 Implementation Details
Top-Hat (Bai and Zhou 2010) Filtering 10.06 75.11 1432 7.14 79.84 1012
Max-Median (Deshpande et al. 1999) 7.00 65.21 59.73 4.17 69.20 55.33
WSLCM (Han et al. 2020) Local Contrast 3.45 72.44 6619 1.16 77.95 5446
TLLCM (Han et al. 2020) 3.31 77.39 6738 1.03 79.09 5899
IPI (Gao et al. 2013) Low Rank 27.92 81.37 16.18 25.67 85.55 11.47
RIPT (Dai and Wu 2017) 14.11 77.55 28.31 11.05 79.08 22.61
ACMNet (Dai et al. 2021b) Deep Learning 59.23 93.27 65.28 70.77 93.08 3.70
ISNet (Zhang et al. 2022) 62.88 92.59 27.92 74.16 97.99 8.35
DNANet (Li et al. 2022) 65.71 91.84 17.61 74.31 98.17 15.97
UIUNet (Wu, Hong, and Chanussot 2022) 65.06 91.16 12.68 72.69 99.08 26.61
HCFNet (Xu et al. 2024) 64.26 92.86 23.91 72.74 98.17 6.21
SCTransNet (Yuan et al. 2024c) 68.64 91.84 11.92 77.09 98.17 15.26
PConv (Yang et al. 2025) 67.08 92.18 11.92 76.25 99.08 6.74
MSHNet (Liu et al. 2024) 67.16 93.88 15.03 74.60 99.08 17.21
@@
EFLNet (Yang et al. 2024) 86.1 29.5 38.0 98.1 37.9 47.1
DFLMF-ISTD (Li et al. 2025) 82.7 38.6 94.2 51.1
PConv (Yang et al. 2025) 86.1 34.5 40.8 96.4 49.9 54.9
@@ 4.3 Ablation Study
Ablation on Different FPNs. To demonstrate the superiority of our NS-FPN, we compare it with representative alternatives including FPN, PANet (Liu et al. 2018), and BiFPN (Tan, Pang, and Le 2020), as shown in Table 4. For segmentation, our method achieves the highest $IoU$ (69.29%) and $P_{d}$ (95.24%), while maintaining acceptable $F_{a}$. For detection, NS-FPN also achieves the best $mAP_{50}$ (86.3%) and $mAP$ (42.1%). These results demonstrate that NS-FPN is an effective plugin designed specifically for the IRSTDS tasks.
@@ 4.4 Comparison with SOTA Methods
Quantitative Comparisons. For segmentation tasks, we select six classical methods and eight state-of-the-art deep learning-based methods for comparisons on the IRSTD-1k and NUAA-SIRST datasets. As presented in Table 5, traditional methods such as Top-Hat (Bai and Zhou 2010) and WSLCM (Han et al. 2020) have limited performance due to their reliance on handcrafted priors. For deep learning-based methods, our method achieves the best performance in all evaluation metrics on both datasets, with 69.29% $IoU$, 95.24% $P_{d}$, and 8.58 $F_{a}$ on IRSTD-1k, and 78.75% $IoU$, 100.0% $P_{d}$, and 1.60 $F_{a}$ on NUAA-SIRST, demonstrating superior performance in pixel-level accuracy and object-level reliability. As for detection tasks, we further compare with four recent IRSTD methods on IRSTD-1k and NUAA-SIRST datasets. Table 5 shows that our approach outperforms previous methods in $mAP_{50}$, $mAP_{75}$ and $mAP$. Specifically, our method achieves the best results on both datasets with 86.3% $mAP_{50}$, 36.9% $mAP_{75}$ and 42.1% $mAP$ on IRSTD-1k, and 97.5% $mAP_{50}$, 61.6% $mAP_{75}$ and 58.0% $mAP$ on NUAA-SIRST. The above results demonstrate that our NS-FPN is highly effective for small target segmentation and detection tasks in infrared scenarios, enabling accurate target localization while suppressing background noise interference.
@@ 4.5 Model Complexity Analysis
Our paper aims to suppress noise interference in feature fusion process by introducing a lightweight yet effective FPN. As shown in Table 4, we can see that NS-FPN has superior segmentation and detection performance compared to other FPNs while ensuring low computational complexity, which verifies the efficiency of NS-FPN. Furthermore, we also analyze the complexity of individual NS-FPN component as shown in Table 6. The parameters and FLOPs for LFP module mainly come from the few convolution operations, which results in similar computational costs with original FPN. Although the SFS module adopts a cross-attention mechanism that increases FLOPs, the overall NS-FPN framework achieves significant performance improvements at an acceptable computational cost compared to original FPN.

### v2

@@ 4.2 Implementation Details
Table 3: LFP at different scale layers are performed on the IRSTD-1k dataset. The best results are highlighted in bold.
@@
Top-Hat (2010) [2] Filtering 10.06 75.11 1432 7.14 79.84 1012
Max-Median (1999) [8] 7.00 65.21 59.73 4.17 69.20 55.33
WSLCM (2020) [12] Local Contrast 3.45 72.44 6619 1.16 77.95 5446
TLLCM (2020) [12] 3.31 77.39 6738 1.03 79.09 5899
IPI (2013) [10] Low Rank 27.92 81.37 16.18 25.67 85.55 11.47
RIPT (2017) [4] 14.11 77.55 28.31 11.05 79.08 22.61
ACMNet (WACV 21) [6] Deep Learning 59.23 93.27 65.28 70.77 93.08 3.70
ISNet (CVPR 22) [33] 62.88 92.59 27.92 74.16 97.99 8.35
DNANet (TIP 22) [13] 65.71 91.84 17.61 74.31 98.17 15.97
UIUNet (TIP 22) [23] 65.06 91.16 12.68 72.69 99.08 26.61
HCFNet (ICME 24) [24] 64.26 92.86 23.91 72.74 98.17 6.21
IRPruneDet (AAAI 24) [35] 64.54 91.74 16.04 75.12 98.61 2.96
IRSAM (ECCV 24) [34] 64.65 90.57 16.61 71.44 92.66 7.53
SCTransNet (TGRS 24) [31] 68.64 91.84 11.92 77.09 98.17 15.26
SFCANet (TAES 25) [17] 66.68 92.89 12.69 78.46 97.24 8.02
PConv (AAAI 25) [26] 67.08 92.18 11.92 76.25 99.08 6.74
MSHNet (CVPR 24) [18] 67.16 93.88 15.03 74.60 99.08 17.21
@@
EFLNet (TGRS 24) [25] 86.1 29.5 38.0 98.1 37.9 47.1
DFLMF-ISTD (IPT 25) [14] 82.7 38.6 94.2 51.1
PConv (AAAI 25) [26] 86.1 34.5 40.8 96.4 49.9 54.9
@@
Figure 6: Heatmaps for the ablation results (IoUIoU, $P_{d}$, $F_{a}$) of $H$ (Rows) and $P$ (Columns) in SFS module.
@@ 4.3 Ablation Study
Ablation of the hyperparameters $H$ and $P$ in SFS. To determine the optimal hyperparameter group of $H$ and $P$ in our SFS modules, we perform ablation studies on IRSTD-1k dataset by varying the number of heads $H\in\{1,2,4,8,16,32\}$ and points $P\in\{3,4,5,6\}$, where the FPN channel dimensions of $64$ should be divisible by $H$. As detailed in Figure 6, increasing $H$ moderately improves performance, but excessive $H$ leads to a performance drop due to insufficient information per head. Similarly, moderate increases in $P$ improve $P_{d}$, while an overly large $P$ introduces more computation and false alarms. Finally, we select the optimal configuration of $H=8$ and $P=4$ for SFS module, where all metrics achieve the best performance.
Ablation on Different FPNs. To demonstrate the superiority of our NS-FPN, we compare it with representative alternatives including FPN [16], PANet [19], BiFPN [22] and HS-FPN [21], as shown in Table 4. For segmentation, our method achieves the highest $IoU$ (69.29%) and $P_{d}$ (95.24%), while maintaining acceptable $F_{a}$. For detection, NS-FPN also achieves the best $mAP_{50}$ (86.3%) and $mAP$ (42.1%). These results demonstrate that NS-FPN is an effective plugin designed specifically for the IRSTDS tasks.
@@ 4.4 Comparison with SOTA Methods
Quantitative Comparisons. For segmentation tasks, we select six classical methods and eleven state-of-the-art deep learning methods, including several frequency domain-based models, for comparisons on the IRSTD-1k and NUAA-SIRST datasets. As presented in Table 5, traditional methods such as Top-Hat [2] and WSLCM [12] have limited performance due to their reliance on handcrafted priors. For deep learning-based methods, our method achieves the best performance in all evaluation metrics on both datasets, with 69.29% $IoU$, 95.24% $P_{d}$, and 8.58 $F_{a}$ on IRSTD-1k, and 78.75% $IoU$, 100.0% $P_{d}$, and 1.60 $F_{a}$ on NUAA-SIRST, demonstrating superior performance in pixel-level accuracy and object-level reliability. As for detection tasks, we further compare with four recent IRSTD methods on IRSTD-1k and NUAA-SIRST datasets. Table 5 shows that our approach outperforms previous methods in $mAP_{50}$, $mAP_{75}$ and $mAP$. Specifically, our method achieves the best results on both datasets with 86.3% $mAP_{50}$, 36.9% $mAP_{75}$ and 42.1% $mAP$ on IRSTD-1k, and 97.5% $mAP_{50}$, 61.6% $mAP_{75}$ and 58.0% $mAP$ on NUAA-SIRST. The above results demonstrate that our NS-FPN is highly effective for small target segmentation and detection tasks in infrared scenarios, enabling accurate target localization while suppressing background noise interference.
@@ 4.5 Model Complexity Analysis
Our paper aims to suppress noise interference in feature fusion process by introducing a lightweight yet effective FPN. As shown in Table 4, we can see that NS-FPN has superior segmentation and detection performance compared to other FPNs while ensuring low computational complexity, which verifies the efficiency of NS-FPN. Furthermore, we also analyze the parameters and computational complexity of individual NS-FPN component as shown in Table 6. The parameters and FLOPs for LFP module mainly come from the few convolution operations, which results in similar computational costs with the original FPN. Although the SFS module adopts a cross-attention mechanism that increases FLOPs, the overall NS-FPN framework achieves significant performance improvements at an acceptable computational cost compared to the original FPN.